OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1

OpenStudy (anonymous):

-7 7 9 -9

OpenStudy (anonymous):

5x+4y = 1 or (1-5x)/4=y 4x+3y = -1 or (-1-4x)/3=y 1-5x = -1-4x 2=x 5(2) + 4y = 1, 4y=-9, y=-9/4 4(2) + 3y = -1 3y=-9, y=-3 The two equations don't seem to have one solution...

OpenStudy (anonymous):

thats what im trying to figure out..

OpenStudy (anonymous):